Blocks 1 and 2(m=2kg,m2=8kg)are hanging with the two blocks connected by a string. Block 1 is pulled upward by an external force, F=150N. Determine the tension in the string that connects the two blocks.
Answer #1

Let the tension between the string be T

writing the force equations we have

150 - 2g - T = 2a ----(1)

T - 8g = 8a ------(2)

adding both equations we have

150 - 10*9.8 = 10a

a = 5.2 m/s^2

T = 8(g + a) = 8(9.8+5.2) = 120 N

so the tension in the string is 120 N